OpenStudy (anonymous):

The balnce B(n+1), in mr. smith's savings account at the end of a year is caculated by the equation B(n+1)=(1.065)Bn, where Bn,= is the balance at the end of the prievious year. Mr. smith made a deposit to open the account 4 years ago. He has not made any withdrawls since. the balance at the end of 2 years was $1701.34. what is the balance at the end of 4 years. Help!!!

OpenStudy (amistre64):

Well, there are a few ways to do this; what have you tried?

OpenStudy (anonymous):

I have tried to use this weird equation they gave me but i dont understand it.

OpenStudy (anonymous):

check on boobears question amistre

OpenStudy (amistre64):

B(n+1)=(1.065)Bn 4 years ago was B(0) 2 years ago B(2) = 1701.34

OpenStudy (amistre64):

the recursion equation, or difference equation, is just a way to feed in a previous amount

OpenStudy (anonymous):

so if it is 4 years later its B(4) ?

OpenStudy (amistre64):

that 1.065 means that he has an interest rate of 6.5% compounded each year

OpenStudy (amistre64):

yes; but I spoose I should have had the starting year as B(1); Balance(in the first year)

OpenStudy (amistre64):

we can work out some simple algebra; but this really just takes the form of a geometric sequence

OpenStudy (anonymous):

? im not in geometry

OpenStudy (amistre64):

B(2) = 1701.34 B(3) = (1.065) * B(2); B(3) = (1.065)(1701.34) right?

OpenStudy (amistre64):

lol.... geometric just means that it is multiplied each time instead of added

OpenStudy (anonymous):

mmmm.... okay what steps would you take to solve this

OpenStudy (amistre64):

B(4) = (1.065) * B(3) ; and we just determined B(3) so lets fill it in

OpenStudy (amistre64):

B(4) = (1.065) * (1.065) * (1701.34)

OpenStudy (amistre64):

B(4) = abt. 1929.70....

OpenStudy (anonymous):

thats one of the answers so i am going say that is right

OpenStudy (anonymous):

thankS

OpenStudy (amistre64):

yay!! :) but do you understand better how I got to it?

OpenStudy (anonymous):

Some what

OpenStudy (amistre64):

if you practice, youll get it :) good luck